# Affective Computing in the Era of Large Language Models: A Survey from the NLP Perspective

### v1

@@ 2.1.1 Sentiment Analysis
AU tasks aim to enhance the model’s ability to recognize and understand various emotions, with Sentiment Analysis (SA) being the core task. The most fundamental task within SA is the Polarity Classification (PC) tasks [64], which generally requires determining whether the sentiment of a given text is negative, positive, or neutral. However, since PC tasks only divide emotions into two or three categories, more detailed Emotional Classification (EC) tasks [65, 66] with more labels have emerged to distinguish emotions better.
Another critical task is Aspect-Based Sentiment Analysis (ABSA) [67, 68], which differs from PC and EC by focusing on sentiment analysis for more specific aspects or attributes. To achieve more fine-grained sentiment analysis, ABSA involves a series of subtasks. The foundation lies in identifying specific aspects or attributes from the text, a process known as Aspect Term Extraction (ATE) [69, 70]. Once these specific aspects are identified, the next step is to perform sentiment analysis for each aspect, known as Aspect-Oriented Sentiment Classification (ASC) [71, 72]. This task aims to determine the user’s sentiment towards each aspect. To further enhance the efficiency and accuracy of the analysis, researchers have proposed Joint Aspect-Sentiment Analysis (JASA) [73]. This method combines aspect extraction and sentiment classification through joint modelling, better capturing the relationship between aspects and sentiments, and enhancing overall performance. Building on this, the Aspect-Category-Opinion-Sentiment (ACOS) task [74] further expands the dimensions of analysis. ACOS considers not only aspects and sentiments but also categories and specific opinions.
In addition to more fine-grained research on emotion labels and aspects, SA tasks also explore the quantification of emotion intensity, known as Emotional Intensity Detection (EID) [75, 76]. EID not only identifies the type of emotion but also assesses its intensity. Furthermore, considering that human emotions are not always explicitly expressed in the text, researchers have proposed the task of Implicit Sentiment Analysis (ISA) [77, 78]. ISA aims to identify and analyze the implicit sentiments in text that are not conveyed through explicit emotional vocabulary or expressions. Moreover, recognizing that emotions can be expressed through various modalities beyond text, researchers have introduced Multimodal Sentiment Analysis (MSA) tasks [79, 80], which integrate textual, visual, and acoustic information to provide a more comprehensive understanding of sentiment.
From the initial SA to more fine-grained ABSA, EID, and ISA, and further to MSA, SA tasks continuously evolve towards more profound, comprehensive, and accurate directions to better understant and apply human emotions.
With the continuous development of human-computer interaction technologies [81], Emotion Recognition in Conversation (ERC) [65, 82, 83] has become increasingly critical. The more challenging task, Emotion and Intent Joint Understanding in Multimodal Conversation (MC-EIU) is proposed. MC-EIU aims to decode the semantic information manifested in a multimodal conversational history while simultaneously inferring the emotions and intents for the current utterance [84]. The task of ERC requires identifying the emotions expressed in each utterance and tracking the emotional dynamics within the context of the conversation to achieve a comprehensive understanding of these dynamics. In a related but distinct direction, the Emotion Cause Pair Extraction (ECPE) [85] focuses on extracting emotions and their corresponding causes from the text. This task necessitates accurately identifying emotional expressions and pinpointing the specific causes of these emotions.
@@ 2.1.2 Subjective Text Analysis
While significant progress has been made in SA and ECR within AC, researchers are simultaneously exploring the potential of Subjective Text Analysis (STA) to provide even deeper insights into user experience and perspective nuances. This burgeoning field of study enriches by delving into the complexities of user intention and attitude revealed through textual information.
In particular, STA encompasses a variety of tasks, such as Personality Assessment [86], Suicide Tendency Detection [87], Toxicity Detection [^1] https://kaggle.com/competitions/jigsaw-toxic-comment-classification-challenge, Sarcasm Detection [88], Well-being Assessment [89], and Engagement Measurement [^2] https://www.kaggle.com/datasets/thedevastator/social-media-interactions-on-tedtalks-dataset, all of which aim to analyze subjective emotions, opinions, and attitudes within text. Compared to classic SA, these tasks receive less attention and fewer research resources.
@@ 2.2.1 Emotional Dialogue
Affective Generation (AG) aims to enhance a system’s capacity to respond to and engage with various emotions. Within this field, Empathetic Response Generation (ERG) [11] and Emotional Support Conversation (ESC) [12] represent two prominent research areas. ERG tasks focus on training models to comprehend a conversation’s content and emotional nuances, enabling them to generate appropriate and empathetic responses. The emergence of multimodal data and research has further propelled this field, with researchers recognizing the efficacy of multimodal responses in fostering empathy within these models, leading to the development of Multimodal Empathetic Response Generation (MERG) [80, 55]. In contrast, ESC tasks adopt a more targeted approach to provide users with professional emotional support and assistance in navigating emotional challenges. Consequently, evaluating emotional support conversation tasks necessitates considering traditional dialogue evaluation metrics and the effectiveness of the emotional support provided [12, 66].
@@ 2.3 Review Summarization
Review Summarization (RS) is a broader task, and opinion summarization is a core component. Opinion Summarization (OS) [90, 91] is the automated process of condensing multiple opinions on a specific topic, typically from social media and user-generated content, into concise and meaningful summaries. The goal is to help users quickly understand a topic’s overall sentiment and key points without reading through large volumes of text. However, it faces challenges due to the unstructured and noisy nature of social media content and issues like sarcasm and diverse language use.
It is worth noting that research on AG lagged behind AU in the pre-LLM era due to limitations in model capabilities and technology. However, with their exceptional sequence generation abilities, the advent of LLMs has ushered in new possibilities for affect generation, making it a focal point of research in the LLM era [92].
@@ 3.1.1 Sentiment Analysis
Sentiment Analysis (SA), a classic task in the Affective Computing domain, focuses on the automatic identification and categorization of emotions and sentiments expressed through contents from users [93, 94, 79].
To evaluate the performance of LLMs in SA tasks, researchers conduct assessments on multiple datasets and discover that the performance of LLMs is closely related to the complexity of the SA tasks they face. For more straightforward sentiment analysis tasks, such as Polarity Classification (PC), closed-source large language models like GPT-3.5-turbo demonstrate robust zero-shot learning capabilities and can deliver satisfactory results [24, 95, 96, 97, 98]. However, when confronted with complex SA tasks requiring deep semantic understanding or structured emotional information, such as Emotional Classification (EC), Aspected-Based Sentiment Analysis (ABSA), Emotional Recognition in Conversation (ERC), etc., which involves identifying multiple sentiments corresponding to different aspects of a post, or Emotion Cause Analysis, the performance of LLMs still lags behind that of pre-trained language models (PLMs) fine-tuned on specific datasets [95, 99, 100, 101]. This discrepancy may stem from LLMs lacking training data tailored to specific tasks and having insufficient capabilities to process complex semantic information.
Although the zero-shot learning capabilities of open-source LLMs are inferior to those of closed-source models, their superior few-shot learning abilities give them an advantage on datasets with limited training data or imbalanced distributions, outperforming PLMs in such scenarios. However, when data is abundant and label distribution is balanced, PLMs can still perform better due to their fine-tuning advantage on specific datasets [102]. Notably, LLMs show superior generalization capabilities in SA tasks compared to PLMs. By simply altering the prompts, LLMs can adapt to different datasets and tasks without requiring model retraining [100].
However, the performance of LLMs is also affected by prompts and decoding parameters. For example, minor text attacks like synonym substitution can lead to fluctuations in their performance. Additionally, because of the auto-regressive decoding mechanism, parameters such as temperature and Top-P can significantly impact the effectiveness of LLMs in sentiment analysis tasks. Typically, a lower temperature value (below 0.3) and a lower Top-P value are necessary for achieving better results [19, 103]. This suggests that although research shows carefully crafted prompts can greatly improve LLMs’performance, developing optimal prompts and setting effective hyper-parameters for specific tasks and contexts remains an unresolved research challenge.
@@ 3.1.2 Subjective Text Analysis
LLMs demonstrate considerable performance variations across various Subjective Text Analysis (STA) related tasks. For instance, they excel in tasks involving negative emotion recognition, such as well-being assessment and toxicity detection. The superior performance could be due to the focus on safety and value alignment during the Reinforcement Learning from Human Feedback (RLHF) stage of LLM training [104, 19, 99]. However, in tasks requiring the understanding of implicit emotional cues, such as engagement measurement, personality assessment, and sarcasm detection, LLMs exhibit a performance gap compared to PLMs fine-tuned on task-specific datasets. This discrepancy may stem from a need for more specialized training data during the supervised fine-tuning (SFT) phase of LLM development. Optimizing prompts for LLMs can enhance their performance on these specific tasks [19, 99]. Furthermore, leveraging their robust capacity for capturing contextual information, LLMs consistently demonstrate superior performance in processing longer texts compared to shorter texts across various subjective text analysis tasks [99, 24]. LLMs perform slightly worse than SFT models tailored to specific tasks in complex sentiment computing areas like sarcasm detection and subtle sentiment analysis. However, the substantial computational and time resources needed to fine-tune these large language models present a significant hurdle. Consequently, developing LLMs with outperformance and minimal training for specific tasks is an important research priority.
@@ 3.2.1 Emotional Dialogue
Empathetic Response Generation (ERG) and Emotional Support Conversation (ESC) are two prominent tasks in AG that attract significant attention, utilizing the EMPATHETICDIALOGUES [11] and ESConv [12], respectively. Early research suggests that LLMs possess immense potential in emotionally-aware generation tasks. For example, GPT-3.5-turbo has shown competitive performance across multiple datasets [4], while GPT-4 generates empathetic responses in various emotional scenarios. Human evaluations indicate that GPT-4’s responses exceed human performance by an average of 10% in empathy levels [105].
Unlike earlier models, LLMs produce longer and more diverse text [105], introducing new evaluation challenges. Traditional automatic evaluation methods based on overlap metrics need to adequately measure the performance of LLMs in emotional generation tasks, while manual evaluation is costly and burdensome to scale. Developing novel evaluation methods and establishing standardized frameworks to accurately measure LLM-generated affective responses’ quality, empathy, and emotional intelligence is crucial for advancing research in this area.
@@ 3.2.2 Emotional Intelligence
Emotional Intelligence (EI) [106, 107, 108] is the ability to recognize, understand, manage, and use emotions effectively in various contexts. In the era of LLMs, assessing and enhancing EI in LLMs is an emerging challenge.
EI is crucial for enhancing the naturalness and effectiveness of human-computer interaction and is a vital component in achieving Artificial General Intelligence (AGI). To explore whether LLMs possess EI, researchers undertake investigations from various perspectives. Current studies[109, 110] draw upon assessment methods from the field of human psychology, employing tools, like the Levels of Emotional Awareness Scale (LEAS) [111], to evaluate the ability of LLMs to perceive emotions. LLMs can generate relatively appropriate emotional responses, which has been evidenced.
However, using human-centric evaluation systems to measure the EI of LLMs has limitations, including the fact that the scales are designed with humans in mind, not explicitly specifically for LLMs and that the evaluations rely on human experts to conduct the assessments. Researchers have developed evaluation tools designed explicitly for LLMs to address these issue, such as the Situational Evaluation of Complex Emotional Understanding (SECEU)[112]. Moreover, to improve the objectivity and efficiency of evaluations, EQ-bench has been proposed [113]. It removes the need for expert assessments and instead directly evaluates the ability of LLMs to understand complex emotions and social interactions by predicting the intensity of emotional states in dialogue characters. While LLMs have demonstrated remarkable performance on SECEU and EQBench, which focus more on the cognition and understanding of human emotional scenarios, there is still insufficient evidence to suggest that they possess emotional intelligence comparable to humans. To address this gap, researchers propose a new task called Sentiment and Opinion Understanding of Language (SOUL), which aims to comprehensively evaluate the model’s ability to understand complex emotions through two sub-tasks, including Review Comprehension and Justification Generation. However, experimental results reveal a substantial performance gap of up to 27% between LLMs and human performance on the SOUL task[114]. It highlights the considerable room for improvement in the capacity of LLMs to understand the nuances and complexities of human emotions.
@@ 4 Instruction tuning
LLMs show excellent performance on different tasks, such as Question Answering, Sequence Generation, and so on [115]. Instruction Tuning [116] is introduced to adapt LLMs to specific tasks, thereby maintaining the benefits of the original model while improving the performance of LLMs on specialized downstream tasks.
Due to the substantial size of LLMs, which often exceed ten billion parameters, Full Parameter Fine-Tuning (FPFT) of all parameters of the complete model could be more practical. To reduce the cost of model training, a series of Parameter-Efficient Fine-Tuning (PEFT) approaches are proposed, such as LoRA [117], Prefix Tuning [118], and P-Tuning [119], as shown in Figure 2. LoRA reduces the number of parameters that need to be trained during the fine-tuning process by freezing all the original parameters and injecting a pair of low-rank decomposition matrices next to the original weights. Prefix Tuning adds trainable "prefix" tokens related to specific tasks as an implicit learnable prompt to the input sequence of the LLM. It involves inserting trainable free vectors into each layer of the Transformer. Applying PEFT in instruction tuning can effectively reduce the training costs in AC tasks.
@@
Figure 2: Illustrations of the instruction tuning methods. Pink blocks denote trainable modules, blue blocks signify frozen modules, and orange blocks represent optionally added modules. Subfigure 2(a) depicts the FPFT method, the transformer blocks of two distinct model architectures are fine-tuned. Subfigure 2(b) introduces the multi-LoRA method. In this approach, only the added multi-LoRA models are trained. This method allows for the optional incorporation of router modules to manage the weights of each LoRA module. Subfigure 2(c) demonstrates the LoRA method, The traditional LoRA method involves the addition of only two low-rank matrix modules, LoRA_A and LoRA_B. VeRA [120] allows the addition of two trainable vectors, d and b, augmenting the LoRA layer. This approach also shares LoRA_A and LoRA_B’s parameters to every LoRA layer. Moreover, when the rank of LoRA in each layer varies, this method can transition into adaLoRA [121]. Subfigure 2(d) illustrates the soft prompt tuning, where a trainable prefix is appended to the model’s embedding layer, known as p-tuning. Furthermore, scientific prefixes can be optionally added to each layer of the model, with only the added prefixes being trained, a technique referred to as prefix-tuning.

### v2

@@ 2.1.1 Sentiment Analysis
AU focuses on recognizing and interpreting human affect, with Sentiment Analysis (SA) as the core task. The basic form is Polarity Classification (PC) Pang et al. (2002), which labels a text as negative, positive, or neutral. Emotional Classification (EC) Li et al. (2017); Demszky et al. (2020) extends this to finer-grained emotion categories. Aspect-Based Sentiment Analysis (ABSA) Schouten & Frasincar (2015); Zhang et al. (2023g) targets aspect-level opinions. Typical subtasks include Aspect Term Extraction (ATE) Toh & Wang (2014); Li et al. (2018) and Aspect-oriented Sentiment Classification (ASC) Lau et al. (2014, 2018); Joint Aspect–Sentiment Analysis (JASA) Zhuang et al. (2020) models both simultaneously, and ACOS Cai et al. (2021) further incorporates category and opinion dimensions. Beyond label granularity and aspects, Emotional Intensity Detection (EID) Bonnet et al. (2015); Alonso et al. (2015) estimates strength, Implicit Sentiment Analysis (ISA) Tubishat et al. (2018); Wei et al. (2020a) captures sentiment not explicitly expressed, and Multimodal Sentiment Analysis (MSA) Das & Singh (2023); Jung et al. (2022) fuses textual, visual, and acoustic cues.
Conversation-level understanding has become central with advances in human–computer interaction Chen et al. (2017). Emotion Recognition in Conversation (ERC) Li et al. (2017); Poria et al. (2019b, 2018) predicts utterance-level emotions while tracking dynamics in context. Emotion and Intent Joint Understanding in Multimodal Conversation (MC-EIU) jointly infers current-utterance emotion and intent Liu et al. (2024b). Emotion Cause Pair Extraction (ECPE) Xia & Ding (2019a) extracts emotions together with their causes. See Table 1 for the complete taxonomy and representative methods.
@@ 2.1.2 Subjective Text Analysis
Subjective Text Analysis (STA) probes user traits, risks, and attitudes expressed in text, including Personality Assessment Ponce-López et al. (2016), Suicide Tendency Detection Desu et al. (2022), Toxicity Detection [^1] https://kaggle.com/competitions/jigsaw-toxic-comment-classification-challenge, Sarcasm Detection Misra & Arora (2023), Well-being Assessment Rastogi et al. (2022), and Engagement Measurement [^2] https://www.kaggle.com/datasets/thedevastator/social-media-interactions-on-tedtalks-dataset. Compared with classic SA, these tasks are relatively under-explored and under-resourced.
@@ 2.2.1 Emotional Dialogue
Affective Generation (AG) aims to produce emotionally appropriate, empathetic, or supportive content. In Emotional Dialogue, Empathetic Response Generation (ERG) Rashkin et al. (2019) generates context-aware empathetic replies, while Emotional Support Conversation (ESC) Liu et al. (2021) targets goal-directed support. Multimodal variants (e.g., MERG) leverage audio/vision with text to enhance empathy and naturalness Jung et al. (2022); Zhang et al. (2024f).
@@ 2.3 Review Summarization
Review Summarization (RS) is a broader task where Opinion Summarization (OS) Kim et al. (2011); Moussa et al. (2018) condenses many user opinions on a topic—often from social media or UGC—into concise summaries that surface overall sentiment and key points. Challenges arise from noisy, unstructured text and figurative or diverse language use (e.g., sarcasm).
It is worth noting that research on AG lagged behind AU in the pre-LLM era due to limitations in model capabilities and technology. However, with their exceptional sequence generation abilities, the advent of LLMs has ushered in new possibilities for affect generation, making it a focal point of research in the LLM era Park et al. (2023).
@@ 3.1.1 Sentiment Analysis
Sentiment Analysis (SA), a classic task in the Affective Computing domain, focus on the automatic identification and categorization of emotions and sentiments expressed through contents from users Tan et al. (2023b); Rodr´ıguez-Ibánez et al. (2023); Das & Singh (2023).
To evaluate the performance of LLMs in SA tasks, researchers conduct assessments on multiple datasets and discover that the performance of LLMs is closely related to the complexity of the SA tasks they face. For more straightforward sentiment analysis tasks, such as Polarity Classification (PC), closed-source large language models like GPT-3.5-turbo demonstrate robust zero-shot learning capabilities and can deliver satisfactory results Krugmann & Hartmann (2024); Zhang et al. (2023f); Belal et al. (2023); Lossio-Ventura et al. (2024); Alexandersen & Rutlin (2024). However, when confronted with complex SA tasks requiring deep semantic understanding or structured emotional information, such as Emotional Classification (EC), Aspected-Based Sentiment Analysis (ABSA), Emotional Recognition in Conversation (ERC), etc., which involves identifying multiple sentiments corresponding to different aspects of a post, or Emotion Cause Analysis, the performance of LLMs still lags behind that of pre-trained language models (PLMs) fine-tuned on specific datasets Zhang et al. (2023f); Amin et al. (2024); Wang et al. (2024h); Wu et al. (2024a). This discrepancy may stem from LLMs lacking training data tailored to specific tasks and having insufficient capabilities to process complex semantic information.
Although the zero-shot learning capabilities of open-source LLMs are inferior to those of closed-source models, their superior few-shot learning abilities give them an advantage on datasets with limited training data or imbalanced distributions, outperforming PLMs in such scenarios. However, when data is abundant and label distribution is balanced, PLMs can still perform better due to their fine-tuning advantage on specific datasets Zhang et al. (2023d). Notably, LLMs show superior generalization capabilities in SA tasks compared to PLMs. By simply altering the prompts, LLMs can adapt to different datasets and tasks without requiring model retraining Wang et al. (2024h). However, the performance of LLMs is also affected by prompts and decoding parameters. For example, minor text attacks like synonym substitution can lead to fluctuations in their performance. Additionally, because of the auto-regressive decoding mechanism, parameters such as temperature and Top-P can significantly impact the effectiveness of LLMs in sentiment analysis tasks. Typically, a lower temperature value (below 0.3) and a lower Top-P value are necessary for achieving better results Amin et al. (2023a); Amin & Schuller (2024). This suggests that although research shows carefully crafted prompts can greatly improve LLMs’performance, developing optimal prompts and setting effective hyper-parameters for specific tasks and contexts remains an unresolved research challenge.
@@ 3.1.2 Subjective Text Analysis
LLMs demonstrate considerable performance variations across various Subjective Text Analysis (STA) related tasks. For instance, they excel in tasks involving negative emotion recognition, such as well-being assessment and toxicity detection. The superior performance could be due to the focus on safety and value alignment during the Reinforcement Learning from Human Feedback (RLHF) stage of LLM training Zheng et al. (2023a); Amin et al. (2023a, 2024). However, in tasks requiring the understanding of implicit emotional cues, such as engagement measurement, personality assessment, and sarcasm detection, LLMs exhibit a performance gap compared to PLMs fine-tuned on task-specific datasets. This discrepancy may stem from a need for more specialized training data during the supervised fine-tuning (SFT) phase of LLM development. Optimizing prompts for LLMs can enhance their performance on these specific tasks Amin et al. (2023a, 2024). Furthermore, leveraging their robust capacity for capturing contextual information, LLMs consistently demonstrate superior performance in processing longer texts compared to shorter texts across various subjective text analysis tasks Amin et al. (2024); Krugmann & Hartmann (2024). LLMs perform slightly worse than SFT models tailored to specific tasks in complex sentiment computing areas like sarcasm detection and subtle sentiment analysis. However, the substantial computational and time resources needed to fine-tune these large language models present a significant hurdle. Consequently, developing LLMs with outperformance and minimal training for specific tasks is an important research priority.
@@ 3.2.1 Emotional Dialogue
Empathetic Response Generation (ERG) and Emotional Support Conversation (ESC) are two prominent tasks in AG that attract significant attention, utilizing the EMPATHETICDIALOGUES Rashkin et al. (2019) and ESConv Liu et al. (2021), respectively. Early research suggests that LLMs possess immense potential in emotionally-aware generation tasks. For example, GPT-3.5-turbo has shown competitive performance across multiple datasets Zhao et al. (2023a), while GPT-4 generates empathetic responses in various emotional scenarios. Human evaluations indicate that GPT-4’s responses exceed human performance by an average of 10% in empathy levels Welivita & Pu (2024). Unlike earlier models, LLMs produce longer and more diverse text Welivita & Pu (2024), introducing new evaluation challenges. Traditional automatic evaluation methods based on overlap metrics need to adequately measure the performance of LLMs in emotional generation tasks, while manual evaluation is costly and burdensome to scale. Developing novel evaluation methods and establishing standardized frameworks to accurately measure LLM-generated affective responses’ quality, empathy, and emotional intelligence is crucial for advancing research in this area.
@@ 3.2.2 Emotional Intelligence
Emotional Intelligence (EI) Dulewicz & Higgs (2000); Mayer et al. (2003); Neubauer & Freudenthaler (2005) is the ability to recognize, understand, manage, and use emotions effectively in various contexts. EI is crucial for enhancing the naturalness and effectiveness of human-computer interaction and is a vital component in achieving Artificial General Intelligence (AGI). To explore whether LLMs possess EI, researchers undertake investigations from various perspectives. Current studies Elyoseph et al. (2024); Ratican & Hutson (2023) draw upon assessment methods from the field of human psychology, employing tools, like the Levels of Emotional Awareness Scale (LEAS) Lane et al. (1990), to evaluate the ability of LLMs to perceive emotions. LLMs can generate relatively appropriate emotional responses, which has been evidenced.
However, using human-centric evaluation systems to measure the EI of LLMs has limitations, including the fact that the scales are designed with humans in mind, not explicitly specifically for LLMs and that the evaluations rely on human experts to conduct the assessments. Researchers have developed evaluation tools designed explicitly for LLMs to address these issue, such as the Situational Evaluation of Complex Emotional Understanding (SECEU) Wang et al. (2023e). Moreover, to improve the objectivity and efficiency of evaluations, EQ-bench has been proposed Paech (2024). It removes the need for expert assessments and instead directly evaluates the ability of LLMs to understand complex emotions and social interactions by predicting the intensity of emotional states in dialogue characters. While LLMs have demonstrated remarkable performance on SECEU and EQBench, which focus more on the cognition and understanding of human emotional scenarios, there is still insufficient evidence to suggest that they possess emotional intelligence comparable to humans. To address this gap, researchers propose a new task called Sentiment and Opinion Understanding of Language (SOUL), which aims to comprehensively evaluate the model’s ability to understand complex emotions through two sub-tasks, including Review Comprehension and Justification Generation. However, experimental results reveal a substantial performance gap of up to 27% between LLMs and human performance on the SOUL Deng et al. (2023d). It highlights the considerable room for improvement in the capacity of LLMs to understand the nuances and complexities of human emotions.
@@ 4 Instruction tuning
LLMs show excellent performance on different tasks, such as Question Answering, Sequence Generation, and so on Zhao et al. (2023b). Instruction Tuning Zhang et al. (2023c) is introduced to adapt LLMs to specific tasks, thereby maintaining the benefits of the original model while improving the performance of LLMs on specialized downstream tasks. Due to the substantial size of LLMs, which often exceed ten billion parameters, Full Parameter Fine-Tuning (FPFT) of all parameters of the complete model could be more practical. To reduce the cost of model training, a series of Parameter-Efficient Fine-Tuning (PEFT) approaches are proposed, such as LoRA Devalal & Karthikeyan (2018), Prefix Tuning Li & Liang (2021), and P-Tuning Liu et al. (2022), as shown in Figure 3. LoRA reduces the number of parameters that need to be trained during the fine-tuning process by freezing all the original parameters and injecting a pair of low-rank decomposition matrices next to the original weights. Prefix Tuning adds trainable "prefix" tokens related to specific tasks as an implicit learnable prompt to the input sequence of the LLM, inserting trainable free vectors into each layer of the Transformer. Applying PEFT in instruction tuning can effectively reduce the training costs in AC tasks.
@@
Figure 3: Illustrations of the instruction tuning methods. Pink blocks denote trainable modules, blue blocks signify frozen modules, and orange blocks represent optionally added modules. Subfigure 3(a) depicts the FPFT method, the transformer blocks of two distinct model architectures are fine-tuned. Subfigure 3(b) introduces the multi-LoRA method. In this approach, only the added multi-LoRA models are trained. This method allows for the optional incorporation of router modules to manage the weights of each LoRA module. Subfigure 3(c) demonstrates the LoRA method, The traditional LoRA method involves the addition of only two low-rank matrix modules, LoRA_A and LoRA_B. VeRA Kopiczko et al. (2024) allows the addition of two trainable vectors, d and b, augmenting the LoRA layer. This approach also shares LoRA_A and LoRA_B’s parameters to every LoRA layer. Moreover, when the rank of LoRA in each layer varies, this method can transition into adaLoRA Zhang et al. (2023b). Subfigure 3(d) illustrates the soft prompt tuning, where a trainable prefix is appended to the model’s embedding layer, known as p-tuning. Furthermore, scientific prefixes can be optionally added to each layer of the model, with only the added prefixes being trained, a technique referred to as prefix-tuning.